OpenStudy (anonymous):

a circle has its origin, and (5,-12) is a point on the circle. how long is the radius of the circle ]

OpenStudy (zpupster):

using the theorem of Pythagoras x^2+y^2 = r^2 (5,-12) or (x,y) r=sqrt((5)^2 + (-12)^2)